# From CLIP to DINO: Visual Encoders Shout in Multi-modal Large Language Models

### v1

@@ Abstract
Multi-modal Large Language Models (MLLMs) have made significant strides in expanding the capabilities of Large Language Models (LLMs) through the incorporation of visual perception interfaces. Despite the emergence of exciting applications and the availability of diverse instruction tuning data, existing approaches often rely on CLIP or its variants as the visual branch, and merely extract features from the deep layers. However, these methods lack a comprehensive analysis of the visual encoders in MLLMs. In this paper, we conduct an extensive investigation into the effectiveness of different vision encoders within MLLMs. Our findings reveal that the shallow layer features of CLIP offer particular advantages for fine-grained tasks such as grounding and region understanding. Surprisingly, the vision-only model DINO, which is not pretrained with text-image alignment, demonstrates promising performance as a visual branch within MLLMs. By simply equipping it with an MLP layer for alignment, DINO surpasses CLIP in fine-grained related perception tasks. Building upon these observations, we propose a simple yet effective feature merging strategy, named COMM, that integrates CLIP and DINO with Multi-level features Merging, to enhance the visual capabilities of MLLMs. We evaluate COMM through comprehensive experiments on a wide range of benchmarks, including image captioning, visual question answering, visual grounding, and object hallucination. Experimental results demonstrate the superior performance of COMM compared to existing methods, showcasing its enhanced visual capabilities within MLLMs. Code will be made available at https://github.com/YuchenLiu98/COMM.
@@ 1 Introduction
Large Language Models (LLMs) (OpenAI, 2023a; b; Touvron et al., 2023a; b; Taori et al., 2023; Chiang et al., ) have made significant strides in the domains of language understanding and generation, achieving remarkable progress recently. Through instruction tuning (Wei et al., 2021; Wang et al., 2022), existing LLMs demonstrate their versatility as general-purpose models capable of handling a wide range of tasks. This capability unlocks their potential zero-shot learning ability, enabling seamless task switching guided by instructions. Building upon the promising performance of LLMs, researchers are now motivated to enhance their capabilities by incorporating visual signals as inputs. This extension allows for the generation of textual outputs that are closely related to visual content, opening up exciting possibilities in the realm of vision-language understanding and generation.
@@
To this end, Flamingo (Alayrac et al., 2022) and BLIP2 (Li et al., 2023c) align the powerful LLMs with a frozen visual encoder to understand visual inputs and perform various vision-language tasks. A series of following works, LLaVA (Liu et al., 2023b), InstructBLIP (Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023) and mPLUG-OWL (Ye et al., 2023) further improve the ability to follow human instructions by constructing multi-modal instruction-following datasets for training. However, these methods are built on image-level alignments, which suffer from the limited fine-grained understanding (such as region description (Liu et al., 2017) and reasoning (Zellers et al., 2019)) and severe object hallucination problem (Li et al., 2023d). To this end, GPT4ROI (Zhang et al., 2023b) proposes instruction tuning on region-of-interest and unlocks the region-level multimodal capacities. Kosmos-2 (Peng et al., 2023) and Shikra (Chen et al., 2023) further integrate the grounding abilities into LLMs and unlock the referential ability in dialogue, i.e., enable the user to point to the object or region as input and the model responds with spatial coordinates of bounding boxes. Such grounding capacity can fulfill numerous vision-language tasks, which is a great progress in MLLMs.
@@
Despite a wide variety of exciting methods and applications, most of existing multi-modal LLMs employ CLIP (Radford et al., 2021) or its variants  (Sun et al., 2023) as the visual branch, where the features output from the deep layers (e.g., the penultimate layer) are usually employed as inputs to the language decoders. However, it still lacks analysis that: Whether using the Vanilla CLIP features as visual encoder is the best way for MLLMs? Though the visual encoder of CLIP is apparently well aligned with the word embedding space by image-text contrastive learning, it fails to learn more detailed pixel-level information such as color and positioning due to the global supervision of image captions, which might hinder the fine-grained perception ability in MLLMs. Besides, existing MLLMs have quite unbalanced visual and language encoders (e.g., ViT-Large-300M vs. Vicuna-7B/13B). Since the language models have succeeded in scaling up the model size with progressively powerful language abilities, the short plate of the Buckets Effect for MLLMs lies in the visual models, which fails to perform in-context learning, and suffer from domain gap and limited zero-shot ability. Consequently, it is critical to enhance the visual capabilities for boosting MLLMs.
@@
This paper presents an extensive investigation into different visual encoders for MLLMs. Four typical visual foundation models are considered, i.e., image-text contrastive learning CLIP, image-only contrastive learning DINOv2 (Oquab et al., 2023), masked image modeling MAE (He et al., 2022) and supervised learning DeiT (Touvron et al., 2021). We evaluate the performance on commonly-used vision-language tasks including visual grounding, object hallucination, visual question answering, image captioning and MME benchmark. Our analysis reveals that different layers of features exhibit varying biases towards local and global patterns. Shallow layer features containing low-level detailed information prove beneficial for fine-grained perception tasks such as grounding and positioning ability, while deep layer features are superior at global understanding. To enhance representation, we propose a multi-level feature merging strategy that incorporates both low-level and high-level features. Surprisingly, when equipped with an MLP layer for alignment, the vision-only model DINOv2 shows promise as a visual branch for MLLMs. We attribute this to the fine-grained localization information captured by DINOv2. Conversely, MAE and DeiT perform inferiorly as visual branches for MLLMs. MAE learns limited semantic information, while DeiT’s strong supervised training makes the alignment with the textual space challenging. Based on the above observations, we propose a fusion strategy that integrates CLIP and DINO with Multi-level features Merging), dubbed as COMM, for boosting the visual branches of MLLMs. Experimental results demonstrate clear advantages of the proposed model over existing approaches and highlight the enhanced visual capabilities brought by COMM.
In a nutshell, the contributions of this paper are summarized as follows:
@@ 2 Related Work
Multi-modal Large Language Model. LLMs (Dai et al., 2019; Brown et al., 2020) have garnered significant attention in both academia and industry due to their remarkable understanding and generative abilities. The success of LLMs has motivated researchers to explore the integration of vision into these models, leading to the development of powerful multi-modal LLMs (MLLMs). Flamingo (Alayrac et al., 2022) employs a cross-attention module to extract visual contexts, which are concatenated with text token as input for LLMs. LLaVA (Liu et al., 2023c) and FROMAGe (Koh et al., 2023) leverage the vision encoder of CLIP to extract visual features, which is aligned to text features using a single linear layer and then input to LLMs. Models such as BLIP-2 (Li et al., 2023b), mPLUG-OWL (Ye et al., 2023), MiniGPT-4 (Zhu et al., 2023) and InstructBLIP (Dai et al., 2023) employ Q-former to extract text-aligned visual features for LLMs. Recently, some interesting works extend LLMs to image retrieval (Koh et al., 2023), video understanding (Zhang et al., 2023a), audio (Su et al., 2023), biomedical analysis (Li et al., 2023a), control systems (Driess et al., 2023).
@@
In recent studies, there has been a growing interest in extending MLLMs to improve their fine-grained understanding abilities through region-level image-text alignment. Kosmos-2 (Peng et al., 2023) addresses this by constructing a large-scale dataset of grounded region-text pairs, enabling the integration of grounding abilities into LLMs. GPT4RoI (Zhang et al., 2023b) reformulates the bounding box as a spatial instruction format and extracts visual features based on region-of-interest, facilitating region-level multi-modal understanding. Shikra (Chen et al., 2023) proposes a unified model that handles spatial coordinates to possess referential abilities in dialogue contexts. Additionally, Qwen (Bai et al., 2023) presents a set of MLLMs that demonstrate remarkable performance across various tasks. However, previous works have predominantly focused on extracting visual features solely from the last few layers of the CLIP model, resulting in an emphasis on global image properties. In this study, we draw attention to the fact that features extracted from shallower layers exhibit a stronger focus on localized properties, which we argue can be more potent in comprehending object locations and image details. Additionally, while CLIP primarily learns globally aligned features, advanced vision-alone models such as DINOv2 excel in capturing more fine-grained vision features. We posit that leveraging these fine-grained vision features can effectively enhance the capabilities of MLLMs, as demonstrated in our analysis. To further advance this line of research, we introduce a novel fusion module that expands and enhances the visual branches, thereby aiming to significantly improve the performance of MLLMs.
@@
Large Vision Foundation Model. Recent progresses in training vision foundation models with large-scale image data focus on contrastive learning, masked image modeling and supervised training. For one thing, contrastive learning can be conducted in an image-only or image-text manner. DINOv2 (Oquab et al., 2023) pretrains the image encoder on large curated image data, which shows a superior understanding of object parts and scene geometry across image domains. Image-text contrastive learning as CLIP (Radford et al., 2021) and EVA-CLIP (Sun et al., 2023) employs the natural language as weak supervision to guide the learning of visual features. For another, BEiT (Bao et al., 2021) predicts discrete tokens based on a pre-trained image tokenizer while iBOT (Zhou et al., 2021) proposes an online image tokenizer. MAE (He et al., 2022) proposes a masked autoencoder for reconstructing image pixels. Besides, DeiT III (Touvron et al., 2022) proposes a training recipe to achieve promising performance. Recent MLLMs employ the vision encoder of CLIP/EVA-CLIP without considering the properties of specific visual models. In this paper, we are the first to re-examine the effectiveness of existing visual models in MLLMs and propose a simple yet effective fusion strategy for boosting visual capabilities.
3 Analysis of the Visual Branch in MLLMs
Previous MLLMs usually utilize the vision encoder of CLIP as their visual branch. Typically, these models extract features from the last few layers, such as the penultimate layer, which are then fed into an alignment network. Subsequently, the aligned features are concatenated with text tokens to serve as input for the LLMs. While the image-text pretraining of CLIP aligns well with the language model, it primarily learns image-level features but overlooks the richer pixel-level features due to the constraint of limited fine-grained information in the guided captions. Moreover, the deep-layer features primarily focus on global image properties and inadequately explore the intricate details of local object parts. As depicted in Fig. 1, the visual features extracted from the shallow layers of CLIP and the deep visual features obtained from the visual-only model DINOv2 contain more detailed information regarding local objects, such as shape or texture. Leveraging these detailed features may enhance the MLLMs’ fine-grained perception abilities.
Evaluation Settings. For further analysis, we conduct a series of quantitative experiments using different kinds of visual models, i.e., image-text contrastive learning CLIP, image-only contrastive learning DINOv2, masked image modeling MAE and supervised learning DeiT. In specific, the visual features extracted from different layers of visual models (based on ViT-Large) are aligned using a linear projection layer and then concatenated with text tokens as the input for LLMs (here we use Vicuna-7B (Chiang et al., )). The overall architecture and training process follow Shikra (Chen et al., 2023) but with fewer iterations (9400 iterations, batch size 16 on 4 A800) to save the computation cost. Then, we measure the capability of the trained MLLMs on referring expression comprehension (REC) (Chen et al., 2023), referring expression generation (REG) (Peng et al., 2023) and object hallucination benchmark (POPE) (Li et al., 2023d). Detailed descriptions of these tasks can be referred to Sec. 5.
(a) Average REC accuracy.
(b) Average POPE accuracy.
(c) Average REG CIDEr.
(a) REC acc for CLIP.
(b) POPE acc for CLIP.
(c) REC acc for DINOv2.
(d) POPE acc for DINOv2
@@ 3 Analysis of the Visual Branch in MLLMs
DINOv2 as the Visual Branch of MLLMs. To leverage the rich fine-grained visual information present in DINOv2, but not inherently aligned with text, we employ a non-linear Multi-Layer Perceptron (MLP) module to align the image features with the word embedding space. Fig. 2 demonstrates that the deep-layer features of DINOv2 exhibit superior grounding abilities, as evidenced by higher REC accuracy, and display satisfactory understanding abilities, as indicated by favorable POPE and REG results. Additionally, we explore the efficacy of multi-level feature merging to enhance performance. In contrast to CLIP, the merging of shallow features from DINOv2 leads to a significant performance degradation. Specifically, in Fig. 3(c) and (d), it is evident that Mean(all) performs notably worse than Mean(19-24) in terms of both REC and POPE accuracy, indicating that the shallow representations lack sufficient semantic information. Building upon the LLN-Layerscale approach, the incorporation of the MLP module for a more potent connection between the visual and text spaces demonstrates a clear improvement in performance. Table 1 showcases the substantial performance gains achieved by employing LLN-Layerscale-MLP as Multi-Level Feature Merging (MFM) module across various vision language tasks. Further detailed ablation studies on the MLP module can be found in the appendix.
@@
MAE and DeiT as the Visual Branch of MLLMs. Fig. 2 shows that MAE features achieve acceptable REC accuracy, but suffers large performance drop on POPE and REG evaluation. This is because MAE features lack sufficient semantic information for global or regional understanding. Therefore, MAE is not suitable as the visual branch for MLLMs. DeiT performs even worse than MAE (details in the appendix). We speculate that this is because supervised training is too strong, which learns a specialized visual space that is difficult to align with the word embedding space.
@@ 4 COMM: Combining CLIP and DINO With Multi-level Feature Merging
Architecture Overview. In this section, we introduce the proposed COMM, that integrates CLIP and DINO with Multi-level features Merging to enhance the visual capabilities of MLLMs. The overall framework is illustrated in Fig. 4, COMM is incorporated into a vision-language instruction following model built upon the recent advanced language and vision-language foundation models. Following the input instructions, our model takes vision and language as inputs to generate text responses following the input instructions. Specifically, we adopt the visual encoder of CLIP and DINOv2 (based on ViT-Large) with our proposed fusion strategy as the visual branch, and Vicuna (Chiang et al., ) (7B/13B) as language decoder. The visual encoder is downsampled with rate 14, meaning that an image with resolution $H\times W$ will be represented by a sequence of $\frac{H}{14}\times\frac{W}{14}$ tokens. The fused token features are projected using a linear layer and then concatenated with the instruction tokens as inputs to the language decoder, which is a generic interface to unify various vision-language tasks as text generation task.
Specifically, denote the visual encoder of CLIP and DINOv2 (ViT Large used) as $f_{1}$ and $f_{2}$, respectively. Given an input image $x$, we extract the patch token features output by all layers of CLIP as $f_{1}(x)=[v_{1}^{1},...,v_{1}^{i},...,v_{1}^{24}]$, where $v_{1}^{i}\in\mathbf{R}^{N\times D}$, $N$ is the number of patch tokens and $D$ is the embedding dimension. The features output by the deep layers of DINOv2 are $f_{2}(x)=[v_{2}^{19},...,v_{2}^{i},...,v_{2}^{24}]$. Then we concatenate the features output by these two models as $\mathbf{v}=[v_{1}^{1},...,v_{1}^{24},v_{2}^{19},...,v_{2}^{24}]$. A linear-layernorm module is leveraged to align the feature space between different layers’ features and layerscale is employed for merging the multiple layer features as,

### v3

@@ Abstract
Multi-modal Large Language Models (MLLMs) have made significant strides in expanding the capabilities of Large Language Models (LLMs) through the incorporation of visual perception interfaces. Despite the emergence of exciting applications and the availability of diverse instruction tuning data, existing approaches often rely on CLIP or its variants as the visual branch, and merely extract features from the deep layers. However, these methods lack a comprehensive analysis of the visual encoders in MLLMs. In this paper, we conduct an extensive investigation into the effectiveness of different vision encoders within MLLMs. Our findings reveal that the shallow layer features of CLIP offer particular advantages for fine-grained tasks such as grounding and region understanding. Surprisingly, the vision-only model DINO, which is not pretrained with text-image alignment, demonstrates promising performance as a visual branch within MLLMs. By simply equipping it with an MLP layer for alignment, DINO surpasses CLIP in fine-grained related perception tasks. Building upon these observations, we propose a simple yet effective feature merging strategy, named COMM, that integrates CLIP and DINO with Multi-level features Merging, to enhance the visual capabilities of MLLMs. We evaluate COMM through comprehensive experiments on a wide range of benchmarks, including image captioning, visual question answering, visual grounding, and object hallucination. Experimental results demonstrate the superior performance of COMM compared to existing methods, showcasing its enhanced visual capabilities within MLLMs.
@@ 1 Introduction
Large Language Models (LLMs) [31, 32, 42, 43, 39, 10] have made significant strides in the domains of language understanding and generation, achieving remarkable progress recently. Through instruction tuning [45, 44], existing LLMs demonstrate their versatility as general-purpose models capable of handling a wide range of tasks. This capability unlocks their potential zero-shot learning ability, enabling seamless task switching guided by instructions. Building upon the promising performance of LLMs, researchers are now motivated to enhance their capabilities by incorporating visual signals as inputs. This extension allows the generation of textual outputs that are closely related to visual content, opening up exciting possibilities in the realm of vision-language understanding.
@@
To this end, Flamingo [1] and BLIP2 [19] align the powerful LLMs with a frozen visual encoder to understand visual inputs and perform various vision-language tasks. A series of following works, LLaVA [23], InstructBLIP [11], MiniGPT-4 [53] and mPLUG-OWL [46] further improve the ability to follow human instructions by constructing multi-modal instruction-following datasets for training. However, these methods are built on image-level alignments, which suffer from the limited fine-grained understanding (such as region description [25] and reasoning [49]) and severe object hallucination problem [21]. To this end, GPT4ROI [51] proposes instruction tuning on region-of-interest and unlocks the region-level multimodal capacities. Kosmos-2 [34] and Shikra [8] further integrate the grounding abilities into LLMs and unlock the referential ability in dialogue, i.e., enable the user to point to the object or region as input and the model responds with spatial coordinates of bounding boxes. Such grounding capacity can fulfill numerous vision-language tasks, which is a great progress in MLLMs.
@@
Despite a wide variety of exciting methods and applications, most of existing multi-modal LLMs employ CLIP [36] or its variants  [38] as the visual branch, where the features output from the deep layers (e.g., the penultimate layer) are usually employed as inputs to the language decoders. However, it still lacks analysis that: Whether using the Vanilla CLIP features as visual encoder is the best way for MLLMs? Though the visual encoder of CLIP is apparently well aligned with the word embedding space by image-text contrastive learning, it fails to learn more detailed pixel-level information such as color and positioning due to the global supervision of image captions, which might hinder the fine-grained perception ability in MLLMs. Besides, existing MLLMs have quite unbalanced visual and language encoders (e.g., ViT-Large-300M vs. Vicuna-7B/13B). Since the language models have succeeded in scaling up the model size with progressively powerful language abilities, the short plate of the Buckets Effect for MLLMs lies in the visual models, which fails to demonstrate emerging capabilities, and suffer from domain gap and limited zero-shot ability. Consequently, it is critical to enhance the visual capabilities for boosting MLLMs.
@@
This paper presents an extensive investigation into different visual encoders for MLLMs. Four typical visual foundation models are considered, i.e., image-text contrastive learning CLIP, image-only contrastive learning DINOv2 [33], masked image modeling MAE [14] and supervised learning DeiT [40]. We evaluate the performance on commonly-used vision-language tasks including visual grounding, object hallucination, visual question answering, image captioning and MME benchmark. Our analysis reveals that different layers of features exhibit varying biases towards local and global patterns. Shallow layer features containing low-level detailed information prove beneficial for fine-grained perception tasks such as grounding and positioning ability, while deep layer features are superior at global understanding. To enhance representation, we propose a multi-level feature merging strategy that incorporates both low-level and high-level features. Surprisingly, when equipped with an MLP layer for alignment, the vision-only model DINOv2 shows promise as a visual branch for MLLMs. We attribute this to the fine-grained localization information captured by DINOv2. Conversely, MAE and DeiT perform inferiorly as visual branches for MLLMs. MAE learns limited semantic information, while DeiT’s strong supervised training makes the alignment with the textual space challenging. Based on the above observations, we propose a fusion strategy that integrates CLIP and DINO with Multi-level features Merging), dubbed as COMM, for boosting the visual branches of MLLMs. Experimental results demonstrate clear advantages of the proposed model over existing approaches and highlight the enhanced visual capabilities brought by COMM. In a nutshell, the contributions of this paper are summarized as follows:
@@ 2 Related Work
Multi-modal Large Language Model. LLMs [12, 5] have garnered significant attention in both academia and industry due to their remarkable understanding and generative abilities. The success of LLMs has motivated researchers to explore the integration of vision into these models, leading to the development of powerful multi-modal LLMs (MLLMs). Flamingo [1] employs a cross-attention module to extract visual contexts, which are concatenated with text token as input for LLMs. LLaVA [24] and FROMAGe [16] leverage the vision encoder of CLIP to extract visual features, which is aligned to text features using a single linear layer and then input to LLMs. Models such as BLIP-2 [20], mPLUG-OWL [46], MiniGPT-4 [53] and InstructBLIP [11] employ Q-former to extract text-aligned visual features for LLMs. Recently, some interesting works extend LLMs to image retrieval [16], video understanding [50], audio [37], biomedical analysis [18], control systems [13].
@@
In recent studies, there has been a growing interest in extending MLLMs to improve their fine-grained understanding abilities through region-level image-text alignment. Kosmos-2 [34] addresses this by constructing a large-scale dataset of grounded region-text pairs, enabling the integration of grounding abilities into LLMs. GPT4RoI [51] reformulates the bounding box as a spatial instruction format and extracts visual features based on region-of-interest, facilitating region-level multi-modal understanding. Shikra [8] proposes a unified model that handles spatial coordinates to possess referential abilities in dialogue contexts. Ferret [47] and ViP-LLaVA [7] further extends with a broader range of free-form shapes for referring, including points, boxes, sketches and scribbles. Additionally, Qwen [3] presents a set of MLLMs that demonstrate remarkable performance across various tasks. However, previous works have predominantly focused on extracting visual features solely from the last few layers of the CLIP model, resulting in an emphasis on global image properties. In this study, we draw attention to the fact that features extracted from shallower layers exhibit a stronger focus on localized properties, which we argue can be more potent in comprehending object locations and image details. Additionally, while CLIP primarily learns globally aligned features, advanced vision-alone models such as DINOv2 excel in capturing more fine-grained vision features. We posit that leveraging these fine-grained vision features can effectively enhance the capabilities of MLLMs, as demonstrated in our analysis. To further advance this line of research, we introduce a novel fusion module that expands and enhances the visual branches, thereby aiming to significantly improve the performance of MLLMs.
Large Vision Foundation Model. Recent progresses in training vision foundation models with large-scale image data focus on contrastive learning, masked image modeling and supervised training. For one thing, contrastive learning can be conducted in an image-only or image-text manner. DINOv2 [33] pretrains the image encoder on large curated image data, which shows a superior understanding of object parts and scene geometry across image domains. Image-text contrastive learning as CLIP [36] and EVA-CLIP [38] employs the natural language as weak supervision to guide the learning of visual features. For another, BEiT [4] predicts discrete tokens based on a pre-trained image tokenizer while iBOT [52] proposes an online image tokenizer. MAE [14] proposes a masked autoencoder for reconstructing image pixels. Besides, DeiT III [41] proposes a training recipe to achieve promising performance. Recent MLLMs employ the vision encoder of CLIP/EVA-CLIP without considering the properties of specific visual models. In this paper, we are the first to re-examine the effectiveness of existing visual models in MLLMs and propose a simple yet effective fusion strategy for boosting visual capabilities.
3 Analysis of the Visual Branch in MLLMs
Previous MLLMs [23, 22, 53, 11, 46, 34, 8, 3, 47] usually utilize the vision encoder of CLIP as their visual branch. Typically, these models extract features from the last few layers, such as the penultimate layer, which are then fed into an alignment network. Subsequently, the aligned features are concatenated with text tokens to serve as input for the LLMs. While the image-text pretraining of CLIP aligns well with the language model, it primarily learns image-level features but overlooks the richer pixel-level features due to the constraint of limited fine-grained information in the guided captions. Moreover, the deep-layer features primarily focus on global image properties and inadequately explore the intricate details of local object parts. As depicted in Fig. 1, the visual features extracted from the shallow layers of CLIP and the deep visual features obtained from the visual-only model DINOv2 contain more detailed information regarding local objects, such as shape or texture. Leveraging these detailed features may enhance the MLLMs’ fine-grained perception abilities.
Evaluation Settings. For further analysis, we conduct a series of quantitative experiments using different kinds of visual models, i.e., image-text contrastive learning CLIP, image-only contrastive learning DINOv2, masked image modeling MAE and supervised learning DeiT. In specific, the visual features extracted from different layers of visual models (based on ViT-Large) are aligned using a linear projection layer and then concatenated with text tokens as the input for LLMs (here we use Vicuna-7B [10]). The overall architecture and training process follow Shikra [8] but with fewer iterations (9400 iterations, batch size 16 on 4 A800) to save the computation cost. Then, we measure the capability of the trained MLLMs on referring expression comprehension (REC) [8], referring expression generation (REG) [34] and object hallucination benchmark (POPE) [21]. Detailed descriptions of these tasks can be referred to Sec. 5.
@@ 3 Analysis of the Visual Branch in MLLMs
DINOv2 as the Visual Branch of MLLMs. To leverage the rich fine-grained visual information present in DINOv2, but not inherently aligned with text, we employ a non-linear Multi-Layer Perceptron (MLP) module to align the image features with the word embedding space. Fig. 2 demonstrates that the deep-layer features of DINOv2 exhibit superior grounding abilities, as evidenced by higher REC accuracy, and display satisfactory understanding abilities, as indicated by favorable POPE and REG results. Additionally, we explore the efficacy of multi-level feature merging to enhance performance. In contrast to CLIP, the merging of shallow features from DINOv2 leads to a significant performance degradation. Specifically, in Fig. 3(c) and (d), it is evident that Mean(all) performs notably worse than Mean(19-24) in terms of both REC and POPE accuracy, indicating that the shallow representations lack sufficient semantic information. Building upon the LLN-Layerscale approach, the incorporation of the MLP module for a more potent connection between the visual and text spaces demonstrates a clear improvement in performance. Table 1 showcases the substantial performance gains achieved by employing LLN-Layerscale-MLP as Multi-Level Feature Merging (MFM) module across various vision language tasks. Further detailed ablation studies on the MLP module are in Section 5.5.
@@
MAE and DeiT as the Visual Branch of MLLMs. Fig. 2 shows that MAE features achieve acceptable REC accuracy, but suffers large performance drop on POPE and REG evaluation. This is because MAE features lack sufficient semantic information for global or regional understanding. Therefore, MAE is not suitable as the visual branch for MLLMs. DeiT performs even worse than MAE (details in Section 5.5). We speculate that this is because supervised training is too strong, which learns a specialized visual space that is difficult to align with the word embedding space.
@@ 4 COMM: Combining CLIP and DINO With Multi-level Feature Merging
Architecture Overview. In this section, we introduce the proposed COMM, that integrates CLIP and DINO with Multi-level features Merging to enhance the visual capabilities of MLLMs. The overall framework is illustrated in Fig. 4, COMM is incorporated into a vision-language instruction following model built upon the recent advanced language and vision-language foundation models. Following the input instructions, our model takes vision and language as inputs to generate text responses following the input instructions. Specifically, we adopt the visual encoder of CLIP and DINOv2 (based on ViT-Large) with our proposed fusion strategy as the visual branch, and Vicuna [10] (7B/13B) as language decoder. The visual encoder is downsampled with rate 14, meaning that an image with resolution $H\times W$ will be represented by a sequence of $\frac{H}{14}\times\frac{W}{14}$ tokens. The fused token features are projected using a linear layer and then concatenated with the instruction tokens as inputs to the language decoder, which is a generic interface to unify various vision-language tasks as text generation task.
Specifically, denote the visual encoder of CLIP and DINOv2 (ViT Large used) as $f_{1}$ and $f_{2}$, respectively. Given an input image $x$, we extract the patch token features output by all layers of CLIP as $f_{1}(x)=[v_{1}^{1},...,v_{1}^{i},...,v_{1}^{24}]$, where $v_{1}^{i}\in\mathbf{R}^{N\times D}$, $N$ is the number of patch tokens and $D$ is the embedding dimension. The features output by the deep layers of DINOv2 are $f_{2}(x)=[v_{2}^{19},...,v_{2}^{i},...,v_{2}^{24}]$. Then we concatenate the features output by these two models as $\mathbf{v}=[v_{1}^{1},...,v_{1}^{24},v_{2}^{19},...,v_{2}^{24}]$. A linear-layernorm module is employed to align the feature space between different layers’ features and layerscale is used to merge multiple layer features as